OpenStudy (anonymous):

Foil Method The "foil" method is a way to multiply binomials horizontally.Solve using the FOIL method. Hint: Watch out for any negative signs and follow rules for signed numbers! (x+4)(x-12)

OpenStudy (anonymous):

= x^2 - 12x + 4x - 48 = x^2 - 8x - 48 the foil method is a memory aid but it is just as easy to remember to first multiply the second bracket by first term in first bracket then by second term in first bracket - the distributive law for multiplication

OpenStudy (anonymous):

so its x^2-8x-48?